What is the half life of Uranium (atomic number=92, atomic mass=238) if 1µCi of radioactivity requires approximately 7.58×10^21 number of 238^U nuclei?
1
Expert's answer
2020-06-12T11:18:03-0400

Number of nuclei is "N=7.58*10^{21}"

Activity of the sample is "R=1 \mu C=3.7×10^{4} decays/s"

The disintegration constant is

"\lambda=\frac{R}{N}=\frac{3.7*10^{4}}{7.58*10^{21}}=4.88*10^{-18} /s"

The half life of the sample is

"T_{1/2}= \frac{ln2}{{4.88*10^{-18}}/s} = 1.42*10^{17} s"